# Improving Code Performance Using LLMs in Zero-Shot: RAPGen

### v2

@@ VI Threats To Validity
In this work, we focus only on the subset of performance issues caused by misuse of common .NET APIs. However, it’s possible to have performance problems that are due to non-API related reasons [Garg et al.(2022)], such as unnecessary computation in a loop that could be hoisted out or excessive allocations of a given type, etc. We also limit our scope to performance issues that can be fixed by changes to a single file and class. However, it’s possible for fixes to span multiple classes, files or folders. This would require a much larger context window as we would need to pass code from multiple classes and the model would then decide which need to be modified in order to fix the issue. A concrete example of this could be a fix that requires making changes to a class that’s used in multiple locations in a given code-base. The fix may require changes to not only the problematic class, but the usages as well. We plan to explore this in future work.
Additionally, we only explore fixing performance issues and assume that there already exists a way to identify expensive parts of a code-base, down to the exact line number. In .NET, this is indeed possible. One could use a profiler [Gayhardt(2019)] and .NET’s Source Link [sou(2018)] technology, which adds metadata containing line number, commit, branch, etc. to a given profiler trace. We have not yet explored this for other languages like Java, python, etc., but we assume that similar technologies exist there as well. We leave this exploration to future work.
Finally, we were unable to run the code fixes suggested to the benchmark dataset. This is because many of the GitHub projects these examples were drawn from were not buildable, for variety of reasons such as, requiring proprietary packages to build, custom build steps or simply compiler errors. However, since our human evaluation showed that the fixes were equivalent to those proposed by the developer, we believe that our changes would result in performance gain as well. In practice, one could add an additional step to run benchmark tests to verify the change improves performance. Furthermore, the benchmark test could itself be generated using an LLM. We leave this exploration to future work. There already exists work that has already explored generating unit tests using these models  [Tufano et al.(2020)].
@@ VII-A Automated Program Repair
Traditional automated program repair (APR) techniques usually require test suits or logic assertions as a specification of correctness. For example, GenProg [Weimer et al.(2009)] uses genetic programming to generate code fixes that pass a supplementary suite of test cases. Pattern-based Automatic program Repair (PAR) [Kim et al.(2013)] leverages manually created fix templates based on patterns in existing developer patches to guide the genetic search. Similarly, CapGen [Wen et al.(2018)] attempts to generate patches at finer granularity (AST node level) using context-aware prioritization of genetic mutation operations, and VarFix [Wong et al.(2021)] further extends GenProg by using variational execution [Austin and Flanagan(2012)]. One of the major challenges with these APR approaches is overfitting [Ye et al.(2021)] and high maintenance cost. They also usually require a significant amount of effort on modeling program semantics. To fill this gap, our approach leverages prompt engineering on LLMs and therefore fixes a wide-range of performance bugs with low training and maintenance cost.
@@ VII-B Neural Code Fix Generation
Recent advancements in deep learning, which automatically learn patterns from unstructured data, have driven a shift in APR techniques away from genetic programming and towards learning-based methods. Machine learning-based APR approaches leverage generative deep learning models that directly output patched versions of faulty code. For example, [Tufano et al.(2018)] uses an RNN encoder-decoder model to generate fixes for bugs on GitHub projects. DeepDebug [Drain et al.(2021)] trained a backtranslation model to generate synthetic buggy versions of code to fine-tune a bug patch model. Stack traces for these generated bugs, as well as supplementary context through code skeletons, are used to finetune a bug patch generation model. Similarly, VRepair [Chen et al.(2022)] trains a vanilla Transformer to automatically repair security vulnerabilities using transfer learning. The model is initially pretrained on bug-fix pairs from GitHub and fine-tuned on security vulnerability-fix pairs. VulRepair [Fu et al.(2022)] builds on VRepair by replacing the vanilla Transformer with a T5 [Raffel et al.(2020)] architecture with BPE tokenization. TFix [Berabi et al.(2021)] is a transformer-based model that fixes bugs found with a static analyzer. More recently, studies have shown that Codex [Chen et al.(2021)] can also generate patches to vulnerable code snippets [Pearce et al.(2021a)]. Similarly, [Prenner et al.(2022)] explored the viability of using Codex in a zero-shot setting to fix bugs in the QuixBugs benchmark [Lin et al.(2017)]. In this setting, Codex was indeed able to generate bug fixes, although fix performance was heavily dependent on the prompts fed to the model, which were created manually. While the majority of the work is focused on general bug fixing or vulnerability fixes, our work uses prompt engineering for fixing performance bugs specifically.
@@ VII-C Prompting and Prompt Engineering
With the rise of LLMs prompt engineering  [Liu et al.(2023)] has emerged as a process to come up with the best prompt for a downstream task. A high quality prompt is often necessary to elicit the correct response from a language model. However, coming up with the optimal prompt is no trivial task [Jiang et al.(2020)]. Prior work have explored various prompt engineering approaches including manually engineering suitable prompts [Petroni et al.(2019), Brown et al.(2020)], few-shot prompting, and optimizing the ordering of prompt examples [Lu et al.(2021)] among others. Similarly, continuous prompts have emerged as an automated prompt learning approach  [Li and Liang(2021), Liu et al.(2021), Lester et al.(2021)]. Continuous prompts directly prompt the embedding space of the model and usually require labeled data. Our work uniquely contributes to prior work by introducing a new prompt generation method called RAPGen, which leverages a knowledge-base of instructions to guide the generation of the prompt.

### v4

@@ VI Threats To Validity
In this work, we focus only on the subset of performance issues caused by misuse of common .NET APIs. However, it’s possible to have performance problems that are due to non-API related reasons [12], such as unnecessary computation in a loop that could be hoisted out or excessive allocations of a given type, etc. We also limit our scope to performance issues that can be fixed by changes to a single file and class. However, it’s possible for fixes to span multiple classes, files or folders. This would require a much larger context window as we would need to pass code from multiple classes and the model would then decide which need to be modified in order to fix the issue. A concrete example of this could be a fix that requires making changes to a class that’s used in multiple locations in a given code-base. The fix may require changes to not only the problematic class, but the usages as well. We plan to explore this in future work.
Additionally, we only explore fixing performance issues and assume that there already exists a way to identify expensive parts of a code-base, down to the exact line number. In .NET, this is indeed possible. One could use a profiler [18] and .NET’s Source Link [27] technology, which adds metadata containing line number, commit, branch, etc. to a given profiler trace. We have not yet explored this for other languages like Java, python, etc., but we assume that similar technologies exist there as well. We leave this exploration to future work.
Finally, we were unable to run the code fixes suggested to the benchmark dataset. This is because many of the GitHub projects these examples were drawn from were not buildable, for variety of reasons such as, requiring proprietary packages to build, custom build steps or simply compiler errors. However, since our human evaluation showed that the fixes were equivalent to those proposed by the developer, we believe that our changes would result in performance gain as well. In practice, one could add an additional step to run benchmark tests to verify the change improves performance. Furthermore, the benchmark test could itself be generated using an LLM. We leave this exploration to future work. There already exists work that has already explored generating unit tests using these models  [28].
@@ VII-A Automated Program Repair
Traditional automated program repair (APR) techniques usually require test suits or logic assertions as a specification of correctness. For example, GenProg [29] uses genetic programming to generate code fixes that pass a supplementary suite of test cases. Pattern-based Automatic program Repair (PAR) [30] leverages manually created fix templates based on patterns in existing developer patches to guide the genetic search. Similarly, CapGen [31] attempts to generate patches at finer granularity (AST node level) using context-aware prioritization of genetic mutation operations, and VarFix [32] further extends GenProg by using variational execution [33]. One of the major challenges with these APR approaches is overfitting [34] and high maintenance cost. They also usually require a significant amount of effort on modeling program semantics. To fill this gap, our approach leverages prompt engineering on LLMs and therefore fixes a wide-range of performance bugs with low training and maintenance cost.
@@ VII-B Neural Code Fix Generation
Recent advancements in deep learning, which automatically learn patterns from unstructured data, have driven a shift in APR techniques away from genetic programming and towards learning-based methods. Machine learning-based APR approaches leverage generative deep learning models that directly output patched versions of faulty code. For example, [35] uses an RNN encoder-decoder model to generate fixes for bugs on GitHub projects. DeepDebug [36] trained a backtranslation model to generate synthetic buggy versions of code to fine-tune a bug patch model. Stack traces for these generated bugs, as well as supplementary context through code skeletons, are used to finetune a bug patch generation model. Similarly, VRepair [37] trains a vanilla Transformer to automatically repair security vulnerabilities using transfer learning. The model is initially pretrained on bug-fix pairs from GitHub and fine-tuned on security vulnerability-fix pairs. VulRepair [38] builds on VRepair by replacing the vanilla Transformer with a T5 [39] architecture with BPE tokenization. TFix [40] is a transformer-based model that fixes bugs found with a static analyzer. More recently, studies have shown that Codex [13] can also generate patches to vulnerable code snippets [41]. Similarly, [42] explored the viability of using Codex in a zero-shot setting to fix bugs in the QuixBugs benchmark [43]. In this setting, Codex was indeed able to generate bug fixes, although fix performance was heavily dependent on the prompts fed to the model, which were created manually. While the majority of the work is focused on general bug fixing or vulnerability fixes, our work uses prompt engineering for fixing performance bugs specifically.
@@ VII-C Prompting and Prompt Engineering
With the rise of LLMs prompt engineering  [44] has emerged as a process to come up with the best prompt for a downstream task. A high quality prompt is often necessary to elicit the correct response from a language model. However, coming up with the optimal prompt is no trivial task [45]. Prior work have explored various prompt engineering approaches including manually engineering suitable prompts [46, 25], few-shot prompting, and optimizing the ordering of prompt examples [47] among others. Similarly, continuous prompts have emerged as an automated prompt learning approach  [48, 49, 50]. Continuous prompts directly prompt the embedding space of the model and usually require labeled data. Our work uniquely contributes to prior work by introducing a new prompt generation method called RAPGen, which leverages a knowledge-base of instructions to guide the generation of the prompt.